OpenStudy (anonymous):

2/3 of what number is 16?

OpenStudy (anonymous):

2/3*24=16

OpenStudy (anonymous):

how did you figure that out lol cuz i have no idea o.o

OpenStudy (anonymous):

ok.....check this, it says 2/3 of some number (say n) is 16.so form the equation 2n/3=16 and hence u'll get n=24. got now?

OpenStudy (anonymous):

2/3 *x=16 x =16*3/2 x=24

OpenStudy (anonymous):

ohh ty so much ~facepalm~ simple ones always get me